an aeroplane flies along a straight line from A to B with speed V and back again with same speed .A steady wind v is blowing . If AB=l then total time for trip if wind blows along AB 

Grade:12

1 Answers

Vikas TU
7 years ago
From the relative velocity concept:
when going from A to B time taken=>
t1 = l/(V-v)
from B to A while returning,
t2 = l /(V+v)
Total Time = t1 + t2 => l /(V+v) + l /(V-v) = > l/(1/(V-v) + 1/(V+v)) seconds.